OpenStudy (anonymous):

Ken is a tour guide at an art museum. He earns $11 an hour regular pay and receives time-and-a-half for overtime. Last week he worked 40 hours plus overtime. If he made $66.00 in overtime, how many overtime time hours did he work?

OpenStudy (anonymous):

over time is time and a half per hour 11/2 + 11 = 5.5 + 11 = 16.50 per hour (this is overtime pay) 66/16.5 = 4 hrs overtime

OpenStudy (anonymous):

do you understand ?

OpenStudy (anonymous):

Yes. It makes sense now. Thank you! @kelliegirl33